# Linear Noise Approximation Assisted Bayesian Inference on Mechanistic Model of Partially Observed Stochastic Reaction Network

### v1

@@ ABSTRACT
To support mechanism online learning and facilitate digital twin development for biomanufacturing processes, this paper develops an efficient Bayesian inference approach for partially observed enzymatic stochastic reaction network (SRN), a fundamental building block of multi-scale bioprocess mechanistic model. To tackle the critical challenges brought by the nonlinear stochastic differential equations (SDEs)-based mechanistic model with partially observed state and having measurement error, an interpretable Bayesian updating linear noise approximation (LNA) metamodel, incorporating the structure information of the mechanistic model, is proposed to approximate the likelihood of observations. Then, an efficient posterior sampling approach is developed by utilizing the gradients of the derived likelihood to speed up the convergence of MCMC. The empirical study demonstrates that the proposed approach has a promising performance.
@@ 1 INTRODUCTION
Partially observed stochastic reaction network (SRN) modeling the dynamics of a population of interacting species, such as chemical molecules participating in multiple reactions, is the fundamental building block of multi-scale bioprocess mechanistic model characterizing the causal interdependences from molecular- to macro-scope dynamics. It plays a critical role to: (1) facilitate digital twin development and support mechanism learning for biomanufacturing processes; (2) allow us to probe critical latent state based on partially observed information; and (3) serve as a fundamental model for a biofoundry platform [1] that can integrate heterogeneous online and offline data collected from different manufacturing processes and speed up the bioprocess development with much less experiments.
@@
SRN has three key features that make the model inference challenging. First, the continuous-time state transition model, representing the evolution of concentration or number of molecules, is highly nonlinear. At any time, the reaction rates, characterizing the regulation mechanisms of enzymatic reaction network, are a function of random state. We adopt the diffusion approximation in [2] to model the state dynamics with a set of coupled stochastic differential equations (SDEs). In this case, the state transition model has double-stochasticity, making it analytically intractable to obtain the state transition densities at different times and also hard to get the closed form likelihood function of observations. Second, since the state is partially observed, we need to integrate out the unobserved state variables to get the likelihood of observations. Third, the data collected from biomanufacturing processes are heterogeneous and also subject to measurement error.
@@
The model inference of enzymatic SRN has found increasing interest especially in biomanufacturing digital twin development. Even under the situations with the reaction network structure known, that is built on thousand years of biological system mechanisms, the mechanistic model parameters are often unknown. It is necessary to infer these parameters using the observations collected from biomanufacturing processes. Since each batch of production can be expensive, we often have very small amount of experimental observations. Coupled with high inherent stochasticity of biomanufacturing processes, the model uncertainty tends to be high. However, frequentist model uncertainty quantification approaches are built on asymptotic approximation, such as asymptotic normality and bootstrap. Thus, in this paper, we focus on a Bayesian inference on multi-scale mechanistic model, which can support online learning and interpretability.
@@
An enormous volume of literature has been dedicated to Bayesian inference for SRN. As the exact state transition density of the nonlinear multivariate SDE-based mechanistic model is unknown, coupled with another challenge (i.e., the partially observed state), the marginal likelihood integrating out the unobserved state variables is intractable. Thus, many existing works are sampling approaches without the explicit calculations of the likelihood, such as approximate Bayesian computation (ABC) and its variants [3]. But in sampling approaches without using likelihood, the complex structure and high stochasticity of SRN make the simulation generating a large amount of sample paths computationally expensive and the acceptance rates of samplers very low respectively. Therefore, we construct a metamodel to approximate the state transition densities of the SDEs, obtain a likelihood approximation, and utilize it to speed up Bayesian inference.
@@
Gaussian Process (GP) is a natural choice to be used as a metamodel. [4] and [5] use GPs as priors for nonparametric estimation of the drift and diffusion terms of SDEs without an exact knowledge of their functional forms. In this paper, we suppose the structure of SDE-based mechanistic model is known. To completely exploit such structure information and improve the interpretability of constructed metamodel, we refer to the deterministic ordinary differential equation (ODE)-based dynamic system inference. In particular, [6] specifies a GP prior over the solution to the ODE, and restricts the GP on a manifold that satisfies the ODE system, to address the incompatibility between the metamodel and the mechanistic model. And an alternative to GP under SDE-based model is linear noise approximation (LNA). The LNA was originally proposed to approximate the solution of the chemical master equation (CME) [7], and it can be derived in a number of ways. For instance, [8] and [9] follow the idea of an asymptotic system size expansion, and derive the LNA by approximating the CME through a Taylor expansion. Since the solution of SDE itself is a random variable, it is difficult to extend the GP approach developed in [6] to the SDE-based model inference. Instead, following [10], we specify the derived LNA as a prior to the solution of the SDE, through which we take full advantage of the structure information provided by the SDE model without the time-consuming numerical integration.
@@
In specific, we propose a LNA assisted Bayesian inference on the nonlinear multivariate SDE-based mechanistic model with partially observed state and subject to measurement error. The main contributions are twofold. First, an interpretable Bayesian updating LNA metamodel is developed for likelihood approximation. It provides a coherent way to simultaneously satisfy the SDE model and fit the observed data, allowing us to probe critical latent state based on partially observed information. Second, the proposed MALA procedure utilizes the gradient information from the derived likelihood to speed up MCMC search and more efficiently generate posterior samples. The proposed Bayesian inference for SRN can support online mechanism learning, facilitate digital twin development, and speed up bioprocess design and control.
@@
2 PROBLEM DESCRIPTION
@@
where $\boldsymbol{v}(\boldsymbol{s}(t);\boldsymbol{\theta})=\left(v_{1}(\boldsymbol{s}(t);\boldsymbol{\theta}_{1}),v_{2}(\boldsymbol{s}(t);\boldsymbol{\theta}_{2}),\ldots,v_{K}(\boldsymbol{s}(t);\boldsymbol{\theta}_{K})\right)^{\top}$ is the reaction rate vector. Equation (3) also represents the doubly stochastic property of SRN, that is, both $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$ are functions of the current system state $\boldsymbol{s}(t)$ and characterized by the parameters $\boldsymbol{\theta}$, while $\boldsymbol{s}(t)$ is a random state vector that changes over time and its evolution (i.e., $d\boldsymbol{s}(t)$) is characterized by $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$.
(2) Partially observed state and heterogeneous data collection. The measures of partially observed state variables are often heterogeneous and subject to measurement error. The observations for different observable state components are also asynchronous; see Figure 1(a). In particular, we represent all observation times of state as the time set, denoted by $\boldsymbol{T}=\{t_{0},t_{1},\ldots,t_{H}\}$, where $t_{0}<t_{1}<\cdots<t_{H}$, and the time intervals $\Delta t_{h}=t_{h+1}-t_{h}$ can be variable for $h=0,1,\ldots,H-1$. At each observation time $t_{h}$, we denote the set of observed components’ subscripts of underlying state $\boldsymbol{s}$ by $\boldsymbol{J}_{h}$, i.e., $\boldsymbol{J}_{h}=\{j\in[J]:s_{j}\text{ is observed at time }t_{h}\}$ where $[J]$ represents $\{1,2,\ldots,J\}$, and let $\boldsymbol{J}_{y}=\cup_{h=0}^{H}\boldsymbol{J}_{h}$ be the set of subscripts of the components that can be observed at certain times of experiments. The observations are denoted by $\boldsymbol{y}_{h}(t_{h})\in\mathbb{R}^{M|\boldsymbol{J}_{h}|}$, where $|\boldsymbol{J}_{h}|\leq J$ is the cardinality of $\boldsymbol{J}_{h}$ representing the dimension of observed components of underlying state $\boldsymbol{s}$ at time $t_{h}$, and $M$ is the batch size of experiments. Then, the observations at time $t_{h}$ can be modeled as
@@
Suppose the measurement error follows a multivariate Gaussian distribution $\boldsymbol{\epsilon}_{h}(t_{h})\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{h})$, where $\boldsymbol{\Sigma}_{h}$ is a diagonal matrix with $M$ vectors of $\boldsymbol{\sigma}_{h}$ on the main diagonal, and $\boldsymbol{\sigma}_{h}=\left\{\sigma_{jj}:j\in\boldsymbol{J}_{h}\right\}^{\top}$ is the vector of measurement error level at time $t_{h}$. Further, let $\boldsymbol{\sigma}=\left\{\sigma_{jj}:j\in\boldsymbol{J}_{y}\right\}^{\top}$ be the vector of measurement error level of all observed components. And $\boldsymbol{G}_{h}$ is a $M|\boldsymbol{J}_{h}|$-by-$J$ constant matrix, mapping the entire $J$-dimensional vector of underlying state $\boldsymbol{s}(t_{h})$ into the $M$ batches of $|\boldsymbol{J}_{h}|$-dimensional vector containing only the counterpart of observed components at time $t_{h}$. Notice that the dimension $|\boldsymbol{J}_{h}|$ can change at different observation times accounting for the fact that the measurements of partially observed state are asynchronous.
@@
(3) Key challenges on Bayesian inference and summary of the proposed inference approach. Our focus in this paper is to develop a computationally efficient Bayesian inference approach on unknown model parameters $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{N}$ with emphasis on nonlinear $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$ characterizing the regulation mechanisms of SRN as shown in (3), where $\boldsymbol{\Theta}^{N}\subset\mathbb{R}^{N}$ is the feasible parameter space. The first challenge is partially observed state subject to random measurement error. The observed components of system state $\boldsymbol{s}$ are recorded at limited discrete time points and the observation time points of each observable component may not be synchronized. Moreover, there are often some components of state $\boldsymbol{s}$ unobservable throughout the experiment. To tackle this challenge, we develop an interpretable Bayesian updating LNA metamodel on underlying state $\boldsymbol{s}(t)$ in Section 3 so that we can predict all components of $\boldsymbol{s}(t)$ at any time $t$.
Such a metamodel needs to have capability to characterize the dependence between components of $\boldsymbol{s}(t)$ and handle the doubly stochasticity of SRN. Luckily, we have a SDE model (3) representing the mechanism of state change. This brings us to the second challenge. On the one hand, the nonlinear drift and diffusion terms of the SDE (3) make solving it directly to get the metamodel of $\boldsymbol{s}(t)$ require time-consuming numerical integration methods. On the other hand, the regulation mechanism structure information from the SDE cannot be completely exploited which impacts on interpretability if we choose a black-box metamodel. To take full advantage of the structure information about the state transition provided by the SDE (3), and to avoid the use of numerical integration to solve the SDE (3), we place LNA priors on the dynamics of state $\boldsymbol{s}(t)$ to facilitate inference of model parameters $\boldsymbol{\theta}$. Under the LNA, the underlying process $\{\boldsymbol{s}(t):t\geq 0\}$ follows a multivariate Gaussian distribution, combined with the assumption of linear Gaussian relation between each observation $\boldsymbol{y}_{h}(t_{h})$ and underlying state value $\boldsymbol{s}(t_{h})$ (i.e., Equation (4)), to give a tractable approximation to the likelihood of the observations $\{\boldsymbol{y}_{h}(t_{h})\}_{h=0}^{H}$. The key to our approach is, to avoid a poor approximation to the true distribution of $\boldsymbol{s}(t)$ as $t$ gets large, we reinitialize the LNA for each time interval $(t_{h},t_{h+1}]$ using the derived posterior distribution of $\boldsymbol{s}(t_{h})$ given $\boldsymbol{y}_{h}(t_{h}),\boldsymbol{y}_{h-1}(t_{h-1}),\ldots,\boldsymbol{y}_{0}(t_{0})$; see Figure 1(b).
@@ 3 BAYESIAN UPDATING LINEAR NOISE APPROXIMATION (LNA) METAMODEL
Step 2: For the subsequent observation time points $t_{1},t_{2},\ldots,t_{H}$, we apply the idea of Kalman filter to sequentially update the LNA prior for $\boldsymbol{s}(t_{h+1})$ and calculate the approximate $p\left(\boldsymbol{y}_{h+1}(t_{h+1})|\boldsymbol{y}_{h}(t_{h}),\ldots,\boldsymbol{y}_{0}(t_{0});\boldsymbol{\theta},\boldsymbol{\sigma}\right)$ recursively for $h=0,1,\ldots,H-1$. Specifically, we first reinitialize the initial values of the ODEs (5) and (7) to the posterior mean and covariance of $\boldsymbol{s}(t_{h})$ respectively. That is, set $\bar{\boldsymbol{s}}(t_{h})=\boldsymbol{\alpha}(t_{h})$ and $\boldsymbol{\Psi}(t_{h})=\boldsymbol{\beta}(t_{h})$. We let $\boldsymbol{\varphi}(t_{h})=0$ as $\boldsymbol{\varphi}(t_{k})=0$ for all $k\geq h$ according to the ODE (6). By integrating the ODEs (5) and (7) through time $t_{h}$ to $t_{h+1}$, we obtain $\bar{\boldsymbol{s}}(t_{h+1})$ and $\boldsymbol{\Psi}(t_{h+1})$. In practice we work with their discretized versions, given by the Euler-Maruyama approximation [13],
@@
As $\Delta t_{h}=t_{h+1}-t_{h}$ is often too large to be used as a time step in (13) and (14), we introduce $\Delta z_{h}=\Delta t_{h}/I_{h}$ for some positive integer $I_{h}\geq 1$. By choosing $I_{h}$ to be sufficiently large, we can ensure the discretization error associated with the Euler-Maruyama approximation is arbitrarily small. That is, to compute $\bar{\boldsymbol{s}}(t_{h+1})$ and $\boldsymbol{\Psi}(t_{h+1})$ more accurately, we recursively calculate the following equations for $i=0,1,\ldots,I_{h}-1$,

### v2

@@ Abstract
To support mechanism online learning and facilitate digital twin development for biomanufacturing processes, this paper develops an efficient Bayesian inference approach for partially observed enzymatic stochastic reaction network (SRN), a fundamental building block of multi-scale bioprocess mechanistic model. To tackle the critical challenges brought by the nonlinear stochastic differential equations (SDEs)-based mechanistic model with partially observed state and having measurement errors, an interpretable Bayesian updating linear noise approximation (LNA) metamodel, incorporating the structure information of the mechanistic model, is proposed to approximate the likelihood of observations. Then, an efficient posterior sampling approach is developed by utilizing the gradients of the derived likelihood to speed up the convergence of Markov chain Monte Carlo (MCMC). The empirical study demonstrates that the proposed approach has a promising performance.
Keywords Partially observed stochastic reaction network  $\cdot$ Bayesian inference  $\cdot$ Linear noise approximation  $\cdot$ Metropolis-adjusted Langevin algorithm
@@ 1 INTRODUCTION
Partially observed stochastic reaction network (SRN) modeling the dynamics of a population of interacting species, such as chemical molecules participating in multiple reactions, is the fundamental building block of multi-scale bioprocess mechanistic model characterizing the causal interdependences from molecular- to macro-kinetics. It plays a critical role to: (1) facilitate digital twin development and support mechanism learning for biomanufacturing processes; (2) allow us to probe critical latent state based on partially observed information; and (3) serve as a fundamental model for a biofoundry platform [1] that can integrate heterogeneous online and offline measures collected from different manufacturing processes and speed up the bioprocess development with much less experiments. Model inference on the SRN mechanistic model based on heterogeneous data also helps to strengthen the theoretical foundations of federated learning on bioprocess mechanisms, through which we can train and advance knowledge.
@@
The SRN mechanistic model has three key features that make the model inference challenging. First, the continuous-time state transition model, representing the evolution of concentration or number of molecules, is highly nonlinear. At any time, the reaction rates, characterizing the regulation mechanisms of enzymatic reaction network, are a function of random state. We adopt the diffusion approximation in [2] to model the state dynamics with a set of coupled stochastic differential equations (SDEs). In this case, the state transition model has double-stochasticity, making it analytically intractable to obtain the state transition densities at different times and also hard to get the closed form likelihood of observations. Second, since the state is partially observed, we need to integrate out the unobserved state variables to get the likelihood. Third, the data collected from biomanufacturing processes are heterogeneous and also subject to measurement errors.
@@
The model inference of enzymatic SRN has found increasing interest especially in biomanufacturing digital twin development. Even under the situations with the reaction network structure known, that is built on thousand years of the understanding on biological system mechanisms, the mechanistic model parameters are often unknown. It is necessary to infer these parameters using the observations collected from biomanufacturing processes. Since each batch of production can be expensive, we often have very small amount of experimental observations. Coupled with high stochasticity of biomanufacturing processes, the model uncertainty tends to be high. However, frequentist model uncertainty quantification approaches are built on asymptotic approximation, such as asymptotic normality and bootstrap. Thus, in this paper, we focus on a Bayesian inference on multi-scale mechanistic model, which can support online learning and interpretability.
@@
An enormous volume of literature has been dedicated to Bayesian inference for SRN mechanistic model. As the exact state transition density of the SDEs-based mechanistic model is unknown, coupled with another challenge (i.e., the partially observed state), the marginal likelihood integrating out the unobserved state variables is intractable. Thus, many existing works are sampling approaches without the explicit calculations of the likelihood, such as approximate Bayesian computation (ABC) and its variants [3]. But in sampling approaches without using likelihood, the complex structure and high stochasticity of SRN make the simulation generating a large amount of sample paths computationally expensive and the acceptance rates of samplers very low. Therefore, we construct a metamodel to approximate the state transition densities of the SDEs, obtain a likelihood approximation, and utilize it to speed up Bayesian inference.
@@
Gaussian Process (GP) is often used as a metamodel. The studies [4] and [5] use GPs as priors for nonparametric estimation of the drift and diffusion terms of SDEs without an exact knowledge of their functional forms. In this paper, we suppose the structure of SDEs-based mechanistic model is known. To completely exploit such structure information and improve the interpretability of constructed metamodel, we refer to the deterministic ordinary differential equation (ODE)-based dynamic system inference. In particular, [6] specifies a GP prior over the solution to the ODE, and restricts the GP on a manifold that satisfies the ODE system, to address the incompatibility between the metamodel and the mechanistic model. And an alternative to GP under SDE-based model is linear noise approximation (LNA). The LNA was originally proposed to approximate the solution of the chemical master equation (CME) [7], and it can be derived in a number of ways. For instance, [8] and [9] follow the idea of an asymptotic system size expansion, and derive the LNA by approximating the CME through a Taylor expansion. Since the solution of SDE itself is a random variable, it is difficult to extend the GP approach developed in [6] to the SDE model inference. Instead, following [10], we specify the derived LNA as a prior to the solution of the SDE, through which we take full advantage of the structure information provided by the SDE model without the time-consuming numerical integration.
@@
In specific, we propose a LNA assisted Bayesian inference on the nonlinear multivariate SDE-based mechanistic model with partially observed state and subject to measurement errors. The main contributions are twofold. First, an interpretable Bayesian updating LNA metamodel is developed for likelihood approximation. It provides a coherent way to simultaneously satisfy the SDE model and fit the observed data, allowing us to probe critical latent state based on partially observed information. Second, the proposed MALA procedure utilizes the gradient information from the derived likelihood to speed up MCMC search and more efficiently generate posterior samples. The proposed Bayesian inference for SRN can support online mechanism learning, facilitate digital twin development, and speed up bioprocess design and control.
@@
2 STOCHASTIC REACTION NETWORK (SRN) MODEL AND PROBLEM DESCRIPTION
@@
where $\boldsymbol{v}(\boldsymbol{s}(t);\boldsymbol{\theta})=\left(v_{1}(\boldsymbol{s}(t);\boldsymbol{\theta}_{1}),v_{2}(\boldsymbol{s}(t);\boldsymbol{\theta}_{2}),\ldots,v_{K}(\boldsymbol{s}(t);\boldsymbol{\theta}_{K})\right)^{\top}$ is the reaction rate vector. Equation (3) also represents the doubly stochastic property of SRN, that is, both mean $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and variance $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$ are functions of the current system state $\boldsymbol{s}(t)$ and characterized by the parameters $\boldsymbol{\theta}$, while $\boldsymbol{s}(t)$ is a random state vector that changes over time and its evolution (i.e., $d\boldsymbol{s}(t)$) is characterized by $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$.
(2) Partially observed state and heterogeneous data collection. The measures of partially observed state variables are often heterogeneous and subject to measurement errors. The observations for different observable state components are also asynchronous; see Figure 1(a). In particular, we represent all observation times of state as the time set, denoted by $\boldsymbol{T}=\{t_{0},t_{1},\ldots,t_{H}\}$, where $t_{0}<t_{1}<\cdots<t_{H}$, and the time intervals $\Delta t_{h}=t_{h+1}-t_{h}$ can be variable for $h=0,1,\ldots,H-1$. At each observation time $t_{h}$, we denote the set of observed components’ subscripts of underlying state $\boldsymbol{s}$ by $\boldsymbol{J}_{h}$, i.e., $\boldsymbol{J}_{h}=\{j\in[J]:s_{j}\text{ is observed at time }t_{h}\}$ where $[J]$ represents $\{1,2,\ldots,J\}$, and let $\boldsymbol{J}_{y}=\cup_{h=0}^{H}\boldsymbol{J}_{h}$ be the set of subscripts of the components that can be observed at certain times of experiments. The observations are denoted by $\boldsymbol{y}_{h}(t_{h})\in\mathbb{R}^{M|\boldsymbol{J}_{h}|}$, where $|\boldsymbol{J}_{h}|\leq J$ is the cardinality of $\boldsymbol{J}_{h}$ representing the dimension of observed components of underlying state $\boldsymbol{s}$ at time $t_{h}$, and $M$ is the batch size of experiments. Then, the observations at time $t_{h}$ can be modeled as
@@
Suppose the measurement errors follow a multivariate Gaussian distribution $\boldsymbol{\epsilon}_{h}(t_{h})\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{h})$, where $\boldsymbol{\Sigma}_{h}$ is a diagonal matrix with $M$ vectors of $\boldsymbol{\sigma}_{h}$ on the main diagonal, and $\boldsymbol{\sigma}_{h}=\left\{\sigma_{jj}:j\in\boldsymbol{J}_{h}\right\}^{\top}$ is the vector of measurement error level at time $t_{h}$. Further, let $\boldsymbol{\sigma}=\left\{\sigma_{jj}:j\in\boldsymbol{J}_{y}\right\}^{\top}$ be the vector of measurement error level of all observed components. And $\boldsymbol{G}_{h}$ is a $M|\boldsymbol{J}_{h}|$-by-$J$ constant matrix, mapping the entire $J$-dimensional vector of underlying state $\boldsymbol{s}(t_{h})$ into the $M$ batches of $|\boldsymbol{J}_{h}|$-dimensional vector containing only the counterpart of observed components at time $t_{h}$. Notice that the dimension $|\boldsymbol{J}_{h}|$ can change at different observation times accounting for the fact that the measures of partially observed state are asynchronous.
@@
(3) Key challenges on Bayesian inference and summary of the proposed inference approach. Our focus in this paper is to develop a computationally efficient Bayesian inference approach on unknown model parameters $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{N}$ with emphasis on nonlinear $\boldsymbol{\mu}(\boldsymbol{s}(t);\boldsymbol{\theta})$ and $\boldsymbol{D}(\boldsymbol{s}(t);\boldsymbol{\theta})$ characterizing the regulation mechanisms of SRN as shown in (3), where $\boldsymbol{\Theta}^{N}\subset\mathbb{R}^{N}$ is the feasible parameter space. The first challenge is partially observed state subject to random measurement error. The observed components of system state $\boldsymbol{s}$ are recorded at limited discrete time points and the observation time points of each observable component may not be synchronized; see Figure 1(a). Moreover, there are often some components of state $\boldsymbol{s}$ unobservable. To tackle this challenge, we develop an interpretable Bayesian updating LNA metamodel on underlying state $\boldsymbol{s}(t)$ in Section 3 so that we can predict all components of $\boldsymbol{s}(t)$ at any time $t$.
Such a metamodel needs to have capability to characterize the dependence between components of $\boldsymbol{s}(t)$ and handle the doubly stochasticity of SRN. Luckily, we have a SDE model (3) representing the mechanism of state change. This brings us to the second challenge. On the one hand, the nonlinear drift and diffusion terms of the SDE (3) make solving it directly to get the metamodel of $\boldsymbol{s}(t)$ require time-consuming numerical integration methods. On the other hand, the regulation mechanism structure information from the SDE cannot be completely exploited which impacts on interpretability if we choose a black-box metamodel. To take full advantage of the structure information about the state transition provided by the SDE (3), and to avoid the use of numerical integration to solve these SDEs, we place LNA priors on the dynamics of state $\boldsymbol{s}(t)$ to facilitate inference of model parameters $\boldsymbol{\theta}$. Under the LNA, the underlying process $\{\boldsymbol{s}(t):t\geq 0\}$ follows a multivariate Gaussian distribution, combined with the assumption of linear Gaussian relation between each observation $\boldsymbol{y}_{h}(t_{h})$ and underlying state value $\boldsymbol{s}(t_{h})$ as shown in Equation (4) to give a tractable approximation to the likelihood of the observations $\{\boldsymbol{y}_{h}(t_{h})\}_{h=0}^{H}$. The key to our approach is, to avoid a poor approximation to the true distribution of $\boldsymbol{s}(t)$ as $t$ gets large, we reinitialize the LNA for each time interval $(t_{h},t_{h+1}]$ using the derived posterior distribution of $\boldsymbol{s}(t_{h})$ given $\boldsymbol{y}_{h}(t_{h}),\boldsymbol{y}_{h-1}(t_{h-1}),\ldots,\boldsymbol{y}_{0}(t_{0})$; see Figure 1(b).
@@ 3 BAYESIAN UPDATING LINEAR NOISE APPROXIMATION (LNA) METAMODEL
Step 2: For the subsequent observation time points $t_{1},t_{2},\ldots,t_{H}$, we apply the idea of Kalman filter to sequentially update the LNA prior for $\boldsymbol{s}(t_{h+1})$ and calculate the approximate $p\left(\boldsymbol{y}_{h+1}(t_{h+1})|\boldsymbol{y}_{h}(t_{h}),\ldots,\boldsymbol{y}_{0}(t_{0});\boldsymbol{\theta},\boldsymbol{\sigma}\right)$ recursively for $h=0,1,\ldots,H-1$. Specifically, we first reinitialize the initial values of the ODEs (5) and (7) to the posterior mean and covariance of $\boldsymbol{s}(t_{h})$ respectively. That is, set $\bar{\boldsymbol{s}}(t_{h})=\boldsymbol{\alpha}(t_{h})$ and $\boldsymbol{\Psi}(t_{h})=\boldsymbol{\beta}(t_{h})$. We let $\boldsymbol{\varphi}(t_{h})=0$ as $\boldsymbol{\varphi}(t_{k})=0$ for all $k\geq h$ according to the ODE (6). By integrating the ODEs (5) and (7) through time $t_{h}$ to $t_{h+1}$, we obtain $\bar{\boldsymbol{s}}(t_{h+1})$ and $\boldsymbol{\Psi}(t_{h+1})$. In practice we work with their discretized versions, given by the Euler method,
@@
As $\Delta t_{h}=t_{h+1}-t_{h}$ is often too large to be used as a time step in (13) and (14), we introduce $\Delta z_{h}=\Delta t_{h}/I_{h}$ for some positive integer $I_{h}\geq 1$. By choosing $I_{h}$ to be sufficiently large, we can ensure the discretization error associated with the Euler method is arbitrarily small. That is, to compute $\bar{\boldsymbol{s}}(t_{h+1})$ and $\boldsymbol{\Psi}(t_{h+1})$ more accurately, we recursively calculate the following equations for $i=0,1,\ldots,I_{h}-1$,